Can u identify the slope and y-intercept, and use them to graph the equation?
An airline frequent-flyer plan offers a bonus of 5000 miles to new members plus 1.5 miles every dollar charged on a credit card endorsed by the airline. The linear equation y=1.5x+5000 represents the number of miles earned after charging x dollars on the credit card.
I tried, but my teacher explains it kinda weird. So far i have y=2 and x=1.9
I'm in 7th grade pre-algebra and this is a take home test and I need a passing grade in this class for my parents to trust me. Thanks.

Answer by stanbon(75887)

Can u identify the slope and y-intercept, and use them to graph the equation?
-------
An airline frequent-flyer plan offers a bonus of 5000 miles to new members plus 1.5 miles every dollar charged on a credit card endorsed by the airline.
---------------
The linear equation y=1.5x+5000 represents the number of miles earned after charging x dollars on the credit card.
-----
The slope is 1.5
Why?
The slope is the change in y associated with a change of 1 in x.
Notice:
If x = 0, y = 1.5*0+5000 = 5000
If x = 1, y = 1.5*1+5000 = 5001.5
---
Notice that y increased by 1.5 when x increased by 1.
============================================================
The intercept is 5000
Why?
The y-intercept is the y-value associated with x = 0.
In your equation, if x = 0, y = 1.5*0+5000 = 5000.
==========================================
Cheers,
Stan H.
============